Combining the Rules to Solve Equations
Solve for x and check your result
9x + 7 = 5x - 3

Answer by bucky(2189)

9x + 7 = 5x - 3
.
The goals in this problem will be to:
.
1) collect all the terms containing x on one side of the equation (by using subtraction)
2) collect all the numbers on the other side of the equation (by using subtraction)
3) combine terms on both sides of the equation (by adding them)
4) solve for x by dividing both sides of the equation by the multiplier of x
5) check by substituting the value of x into the original equation to make sure both sides
are equal.
.
Step 1):
.
Get rid of the 5x on the right side by subtracting 5x from both sides to get:
.
9x -5x + 7 = -3
.
Step 2):
.
Get rid of the + 7 on the left side by subtracting +7 from both sides:
.
9x - 5x = -3 - 7
.
Step 3):
.
Combine the terms on both sides and this results in:
.
4x = -10
.
Step 4):
.
Solve for x by dividing both sides by the multiplier of x (multiplier = 4):
.
4x/4 = -10/4
.
and this reduces to:
.
x = -10/4 = -5/2
.
So we have found that x = -5/2
.
Step 5)
.
Check by substituting -5/2 for x in the original given equation:
.
9x + 7 = 5x - 3
.
9*(-5/2) + 7 = 5*(-5/2) - 3
.
Multiply out to get:
.
-45/2 + 7 = -25/2 -3
.
Multiply all terms of this equation (on both sides) by 2 to get rid of the denominators:
.
2*(-45/2) + 2*7 = 2*(-25/2) + 2*(-3)
.
-45 + 14 = -25 - 6
.
-31 = -31
.
Both sides of the equation are equal, so the answer is correct ... x = -5/2 which also
can be written as x = -2.5 or x = -2 1/2.
.
Hope this helps you to understand the problem and the procedures that you use to solve
linear equations of this type.
.